a lookout tower casts a shadow 200 feet long at the same time that the shadow of a 8 foot truck is 16 feet long. find the height of the tower.
Answer by stanbon(75887)

a lookout tower casts a shadow 200 feet long at the same time that the shadow of a 8 foot truck is 16 feet long. find the height of the tower.
---
tower/shadow = truck/shadow
----
t/200 = 8/16
----
t = 200(1/2)
----
t = 100 ft (height of the tower)
===================
Cheers,
Stan H.
===================